Physics question.
2) Four resistors are connected across an 8-V DC b
Answer #1

cz it is parallel combination of resistance

so the voltage at every branch will be same

so voltage at branch 1 (3 ohm and 9 ohm) will be

V1=VxR1/(R1+R2)

V1=8x12/18

V1=5.33

total current through 12 ohm

I1=V1/R1

I1=5.33/12

I1=0.444

from current division rule

current through 9 ohm

I=.444x12/9

I=0.58666666666